OpenStudy (anonymous):

What is the domain and range of the quadratic equation y = -x2 - 14x - 52?

OpenStudy (jdoe0001):

for a parabola equation like \(\bf \Large y = ax^2+bx+c\) you can find its vertex at the coordinates of \(\bf \left(-\cfrac{b}{2a}\quad ,\quad c-\cfrac{b^2}{4a}\right)\) so what would you get for the vertex?

OpenStudy (jdoe0001):

anyhow, the domain, that is, values that "x" can take, are pretty much \(\bf x\in \mathbb{R}\) the range on the other hand, is a bit constrained, once you get the vertex, you'll notice the range

OpenStudy (anonymous):

ahhh! thank you(: